VEGAS Team Meeting

January 22, 2013


CALTECH, UC-B, & Columbia

Dan, Jayanth, Mark, Andrew


Jason, Ray, Marty, Anish, Mark

Meeting Notes:

Next Teleconference: January 28, 2012 .

VEGAS observation preparation Status:

  1. half part of the IF band has poor SNR. why? is it fixed already? how to fix?
    • DONE Resolved and ready for test
  2. the start up problem:
    • DONE Resolved and ready for test
  3. 16 bins of the spectra have poor SNR. (an FPGA design problem).
    • DONE Resolved and ready for test
  4. valon locking problem.
    • DONE Solved by a power cycle - inconsistent status between python GUI and laptop
  5. what else needs to get fixed before we observe?
    • DONE 12 second gaps
      • Resolved and ready for test
  6. what tests should we do before we have enough confidence to start observing?
    • Test Pulsar
      • Ready to go - test 22-Jan PM
        • version with 'broken-up' observation data to later verify the ability to stitch together multiple observations
      • Broadband noise test 23 Jan AM
    • Known Pulsar - with gap to prove obs can be combined
      • Scheduled GBT time 23-Jan PM
      • On-Off continuum obs for sensitivity
    • Six hour data collection test period would be nice - can be done with test pulsar

Action Items:

Other items:

-- MartyBloss - 2013-01-22
Topic attachments
I Attachment Action Size Date Who Comment
VEGAS_Pre_130122.docxdocx VEGAS_Pre_130122.docx manage 13 K 2013-01-22 - 15:07 MartyBloss Test Planning Document for January 22-24, 2013
Topic revision: r1 - 2013-01-22, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback